OpenStudy (anonymous):

Rewrite the expression using rational exponent notation or radical notation.

OpenStudy (anonymous):

\[(\sqrt{y})^7\]

OpenStudy (anonymous):

\[(\sqrt[a]{x})^b= x^{b/a}\]

OpenStudy (anonymous):

\[y^7/^2\]

OpenStudy (anonymous):

Is that it?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem :D